# Adaptive Multi-Order Graph Regularized NMF with Dual Sparsity for Hyperspectral Unmixing

### v1

@@ I Introduction
During the past decades, a large number of methods have emerged to deal with the HU problem under the framework of a linear mixing model (LMM), which can be roughly divided into geometry-based [16, 17], statistics-based [18, 19], regression-based [20, 21], and deep learning-based [22, 23]. Specifically, HU includes two stages: endmember extraction, such as vertex component analysis (VCA) [24] based on the geometric structure, and abundance estimation, such as fully constrained least squares (FCLS) [25] based on least squares. Actually, VCA-FCLS has been broadly used for initialization in the unmixing process. It is worth noting that nonnegative matrix factorization (NMF) [26], as a statistics-based method, has been recognized as a powerful unmixing tool because of its nonnegative property consistent with the fact that image pixel values are nonnegative while allowing the simultaneous estimation of endmembers and their respective abundances. See [27, 28] for recent surveys about NMF.
@@
To enhance the unmixing performance of NMF, researchers have developed a large number of constrained NMF methods and achieved remarkable results. For example, Miao and Qi [29] incorporated a volume constraint based on the convex cone simplex into the standard NMF and proposed MVC-NMF. Qian et al. [30] suggested $\ell_{1/2}$-norm [31] constrained NMF to improve the interpretability, which is denoted as SNMF in our paper without ambiguity. Subsequently, Zhu et al. [32] replaced $\ell_{1/2}$-norm with $\ell_{q}$-norm $(0<q<1)$ to obtain the ability of adaptive selection [33]. Furthermore, Wang et al. [34] considered the sparse structure of local pixels and combined it with the smoothness constraint, which is called spatial group sparsity regularized NMF (SGSNMF). Besides, there are low-rank constrained NMF [35], [36], abundance separation constrained NMF [37], [38], etc. We would like to point out that although constrained NMF methods significantly improve the performance of traditional NMF by incorporating prior information, it is still not robust to noise and outliers [39, 40]. To this end, Févotte et al. [41] proposed robust NMF (RNMF) by introducing additional sparse matrices to filter out sparsely distributed noise. Recently, Zhang et al. [42] constructed group sparsity constrained NMF, which is denoted as IRNMF. In addition to the above constrained NMF methods, Abdolali et al. [43] transformed the minimum-volume NMF problem in the primal space to a maximum-volume problem in the dual space for a more stable global solution, which is called MV-DUAL. In summary, the above-mentioned sparse constrained NMF methods have good unmixing performance, hence we will continue to align this strategy.
@@
Recently, graph learning was incorporated with NMF to capture the low-dimensional features in HSIs [44, 45]. Lu et al. [46] proposed a graph-regularized NMF method by combining manifold regularization with sparse constrained NMF, where the manifold regularization term was validated to strengthen the connection between original images and abundance maps. Yang et al. [47] considered spatial geometric distance and spectral geometric distance as metrics and obtained a space-spectral manifold-regularized NMF model. Wang et al. [48] introduced the hypergraph structure to capture the similarity relationships among spatially neighboring pixels and proposed hypergraph-regularized sparse NMF. Very recently, Qu et al. [49] presented a high-order graph regularized NMF method, called HGNMF-FS. However, these graph learning-based methods have two limitations. On the one hand, they mainly focus on first-order or fixed-order neighborhood relations [50]. This may lead to the loss of key pixel association information when applied to HSIs with complex distributions, thereby limiting the model’s ability to represent global and local structures. On the other hand, their graph construction always relies on the predefined parameters and lacks the ability to adaptively learn graph structures.
@@ II-C Nonnegative Matrix Factorization
[…]
@@ II-D Graph Constrained NMF
$\rm{s.t.}{}{}{}{}$ $\mathbf{A}\geq 0,~{}\mathbf{S}\geq 0,~{}\mathbf{\Lambda}\geq 0,~{}\mathbf{1}_{R}^{\top}\mathbf{S}=\mathbf{1}_{N}^{\top}.$
@@ III-A Spatial and Spectral Graphs
where $\sigma_{s}$ is the spatial scaling parameter that controls the influence of distance between neighboring pixels. In this way, the spatial map can capture the local spatial structure of pixels, while Gaussian weighting can smooth the similarity between neighboring pixels.
@@
Therefore, the corresponding first-order Laplacian matrices of the spatial and spectral maps are defined as
@@ III-B Multi-Order Graph
Although constructing first-order nearest neighbor graphs can achieve a low-dimensional approximation of adjacent pixel representations, such local connectivity constraints often fail to characterize the global information inherent in HSIs, where the similarity between many pixels may extend beyond direct neighbors. In this regard, high-order graphs can be considered to improve the accuracy of HU [52]. Let $\mathbf{W}\in\mathbb{R}^{N\times N}$ denote the first-order graph with $N$ nodes, and the $k$-order graph $\mathbf{W}_{k}$ is defined as
@@
$\mathbf{W}_{m}$ $=\beta_{1}\mathbf{W}_{1}+\beta_{2}\mathbf{W}_{2}+\dots+\beta_{k}\mathbf{W}_{k},$ (12)
$\mathbf{D}_{m}$ $=\beta_{1}\mathbf{D}_{1}+\beta_{2}\mathbf{D}_{2}+\dots+\beta_{k}\mathbf{D}_{k},$
$\mathbf{L}_{m}$ $=\beta_{1}\mathbf{L}_{1}+\beta_{2}\mathbf{L}_{2}+\dots+\beta_{k}\mathbf{L}_{k},$
where $\beta_{1},\dots,\beta_{k}$ are the weight parameters. From both spatial and spectral perspectives, we construct the following spatial and spectral multi-order graph
$\mathbf{W}_{m}$ $=\beta_{1}\mathbf{W}_{1}^{\text{spa}}+\beta_{2}\mathbf{W}_{2}^{\text{spa}}+\dots+\beta_{k}\mathbf{W}_{k}^{\text{spa}}$ (13)
@@
$\mathbf{D}_{m}$ $=\beta_{1}\mathbf{D}_{1}^{\text{spa}}+\beta_{2}\mathbf{D}_{2}^{\text{spa}}+\dots+\beta_{k}\mathbf{D}_{k}^{\text{spa}}$
@@
$\mathbf{L}_{m}$ $=\beta_{1}\mathbf{L}_{1}^{\text{spa}}+\beta_{2}\mathbf{L}_{2}^{\text{spa}}+\dots+\beta_{k}\mathbf{L}_{k}^{\text{spa}}$
@@
$=\text{Tr}(\mathbf{S}\mathbf{L}_{m}\mathbf{S}^{\top}).$
@@ III-C Our Model
s.t. $\mathbf{A}\geq 0,~{}\mathbf{S}\geq 0,~{}\mathbf{1}_{R}^{\top}\mathbf{S}=\mathbf{1}_{N}^{\top},$
@@ IV-A3 Update $\mathbf{E}$
Finally, Algorithm 1 summarizes the overall framework for solving (15), where the paramers of multi-order graphs in updating $\mathbf{S}$ can be adaptively learned through Algorithm 2.
@@
Initialize: $(\mathbf{A}^{0},\mathbf{S}^{0},\mathbf{E}^{0})$
While not converged do
@@
End While
@@ IV-B1 Update $\mathbf{W}_{m}$
$\mathbf{W}_{m}\leftarrow\max\left(0,\sum_{v=1}^{V}\sum_{k=1}^{K}H_{vk}\mathbf{W}_{k}^{v}/\left(2\left(1+\mu\right)\right)\right).$ (30)
@@ IV-B2 Update $\mathbf{H}$
Initialize: $H_{vk}=\frac{1}{2K}$, multi-order graphs via (10)
While not converged do
@@
End While
@@ V Numerical Experiments
This section presents numerical experiments between our proposed MOGNMF and some state-of-the-art HU methods, including VCA-FCLS [24] [25], MVC-NMF [29], SNMF [30], RNMF [41], SGSNMF [34], IRNMF [42], HGNMF-FS [49], and MV-DUAL [43].
@@
Figure 2: Parameter sensitivity analysis with respect to $\lambda$ and $\beta$, where (a) SAD and (b) RMSE.
@@ V-A1 Parameter Settings
In the experiments, all compared methods are tested with 20 iterations to ensure the validity and all parameters are chosen according to their literatures. For our proposed MOGNMF, $\gamma$, $\beta$, and $\lambda$ should be taken into account. According to [30], $\gamma$ is estimated by
@@
where $L$ represents the number of spectral bands, $N$ represents the total number of pixels, and $\mathbf{x}^{l}$ represents the vector corresponding to the $l$-th band in the hyperspectral data $\mathbf{X}\in\mathbb{R}^{L\times N}$. According to Fig. 2, it can be seen that SAD and RMSE exhibit similar variation trends, thus, in subsequent experiments, we set $\beta=1.5$ and $\lambda=0.01$. In Algorithm 2, $\mu$ is empirically set to 0.01 and $\alpha$ is set to 0.1.
TABLE I: SAD values of different methods on the Simu-1 dataset, where the top two values are marked as red and blue, respectively.
  Cases   VCA-FCLS   MVC-NMF   SNMF   RNMF   SGSNMF   IRNMF   HGNMF-FS   MV-DUAL  MOGNMF
SNR=10 0.0968 0.0356 0.0442 0.0627 0.0358 0.0361 0.0438 0.0982 0.0320
SNR=20 0.0821 0.0350 0.0524 0.0610 0.0436 0.0340 0.0331 0.0443 0.0313
SNR=30 0.0321 0.0109 0.0467 0.0508 0.0468 0.0119 0.0319 0.0320 0.0099
SNR=40 0.0122 0.0090 0.0460 0.0149 0.0393 0.0076 0.0267 0.0335 0.0063
Average 0.0558 0.0226 0.0473 0.0473 0.0432 0.0224 0.0338 0.0520 0.0202
TABLE II: RMSE values of different methods on the Simu-1 dataset, where the top two values are marked as red and blue, respectively.
 Cases   VCA-FCLS   MVC-NMF   SNMF   RNMF   SGSNMF   IRNMF   HGNMF-FS   MV-DUAL  MOGNMF
SNR=10 0.0943 0.0966 0.0988 0.0932 0.0944 0.0935 0.1180 0.1517 0.0899
SNR=20 0.0924 0.0937 0.1008 0.1151 0.0941 0.0955 0.0925 0.1276 0.0917
SNR=30 0.0552 0.0310 0.0894 0.0600 0.0599 0.0533 0.0515 0.0475 0.0276
SNR=40 0.0391 0.0319 0.0916 0.0392 0.0881 0.0376 0.0619 0.0462 0.0219
Average 0.0703 0.0633 0.0952 0.0769 0.0841 0.0700 0.0810 0.0933 0.0578

### v2

@@ I Introduction
Based on the spectral mixing mechanism, a large number of methods have been proposed to deal with the HU problem. The linear mixing model (LMM) and the nonlinear mixing model (NLMM) are two representatives. LMM assumes that each pixel is a linear combination of endmember spectra, while NLMM further accounts for multiple scattering and microscopic mixing effects [16]. Typical NLMM methods include the generalized bilinear model [17], the p-order polynomial model [18], and the multilinear mixing model [19]. Due to the inherently nonconvex nature of unmixing problems, nonlinear methods often involve high degrees of freedom, which may lead to the generation of physically meaningless endmembers [20]. Nowadays, LMM has been widely regarded as a reasonable first-order approximation of the physical measurement process due to its strong physical interpretability[21].
Generally speaking, LMM can be divided into geometry-based [22, 23], statistics-based [24, 25], regression-based [26, 27], and deep learning-based [28, 29]. Geometry-based methods like vertex component analysis (VCA) [30] extract endmembers by locating extreme pixels according to the data’s simplex structure, but they usually require pure pixels. Regression-based methods such as fully constrained least squares (FCLS) [31] estimate abundances under nonnegativity and sum-to-one constraints. The combination of VCA and FCLS, known as the VCA-FCLS framework, is widely used as initialization by separately handling endmember extraction and abundance estimation. In contrast, nonnegative matrix factorization (NMF) [32] jointly estimates endmembers and abundances under nonnegativity constraints, aligning with the physical properties of hyperspectral data. Unlike VCA-FCLS, it does not require pure pixels and supports unsupervised learning. Its flexibility and extensibility have made NMF a widely used tool, as highlighted in recent surveys [33, 34].
@@
To enhance the unmixing performance of NMF, researchers have developed a large number of constrained NMF methods and achieved remarkable results. For example, Miao and Qi [35] incorporated a volume constraint based on the convex cone simplex into the standard NMF and proposed MVC-NMF. Qian et al. [36] suggested $\ell_{1/2}$-norm [37] constrained NMF to improve the interpretability, which is denoted as SNMF in our paper without ambiguity. Subsequently, Zhu et al. [38] replaced $\ell_{1/2}$-norm with $\ell_{q}$-norm $(0<q<1)$ to obtain the ability of adaptive selection [39]. Furthermore, Wang et al. [40] considered the sparse structure of local pixels and combined it with the smoothness constraint, which is called spatial group sparsity regularized NMF (SGSNMF). Besides, there are low-rank constrained NMF [41], [42], abundance separation constrained NMF [43], [44], etc. We would like to point out that although constrained NMF methods significantly improve the performance of traditional NMF by incorporating prior information, it is still not robust to noise and outliers [45, 46]. To this end, Févotte et al. [47] proposed robust NMF (RNMF) by introducing additional sparse matrices to filter out sparsely distributed noise. Recently, Zhang et al. [48] constructed group sparse constrained NMF, which is denoted as IRNMF. In addition to the above constrained NMF methods, Abdolali et al. [49] transformed the minimum-volume NMF problem in the primal space to a maximum-volume problem in the dual space for a more stable global solution, which is called MV-DUAL. In summary, the above-mentioned sparse constrained NMF methods have good unmixing performance, hence we will continue to align this strategy.
@@
Recently, graph learning was incorporated with NMF to capture the low-dimensional features in HSIs [50, 51]. Lu et al. [52] proposed a graph-regularized NMF method by combining manifold regularization with sparse constrained NMF, where the manifold regularization term was validated to strengthen the connection between original images and abundance maps. Yang et al. [53] considered spatial geometric distance and spectral geometric distance as metrics and obtained a space-spectral manifold-regularized NMF model. Wang et al. [54] introduced the hypergraph structure to capture the similarity relationships among spatially neighboring pixels and proposed hypergraph-regularized sparse NMF. Very recently, Qu et al. [55] presented a high-order graph regularized NMF method, called HGNMF-FS. However, these graph learning-based methods have two limitations. On the one hand, most methods rely on first-order or fixed-order neighborhoods, resulting in sparse graphs that fail to capture high-order pixel associations [56]. This limitation reduces the model’s ability to represent both global and local structures, such as the similarity between neighboring pixels and the consistency across distant but similar regions [57]. On the other hand, their graph construction typically depends on predefined parameters, lacking the flexibility to adaptively learn graph structures. As shown in Fig. 1, adaptive learning allows the model to discover the optimal graph as a graph regularization constraint, where the complementary nature of different orders significantly enhances the expressiveness of the graph representation.
In addition, deep learning can effectively model complex nonlinear mixtures, thereby capturing hierarchies in hyperspectral data. Tao et al. [58] addressed the strong reliance of existing unmixing methods on initialization, and proposed an abundance-guided attention network called A2SAN. Chen et al. [59] leveraged the Mamba model to achieve efficient hyperspectral unmixing with fewer parameters and lower computational cost. Very recently, Che et al. [60] learned multi-order graphs from different views, thereby capturing high-order relationships between data. Yang et al. [61] embedded graph regularization constraints in each view and adaptively assigned weights to different views. These two methods can effectively extract valuable information from multi-view data.
@@ II-C Nonnegative Matrix Factorization
where the operator $\cdot\ast$ and ././ represent element-wise multiplication and division, respectively.
@@ II-D Graph Constrained NMF
$\rm{s.t.}$ $\mathbf{A}\geq 0,~\mathbf{S}\geq 0,~\mathbf{\Lambda}\geq 0,~\mathbf{1}_{M}^{\top}\mathbf{S}=\mathbf{1}_{N}^{\top}.$
@@ III-A Spatial and Spectral Graphs
where $\sigma_{s}$ is the spatial scaling parameter that controls the influence of the distance between neighboring pixels. In this way, the spatial graph can capture the local spatial structure of pixels, while Gaussian weighting can smooth the similarity between neighboring pixels.
@@
Therefore, the corresponding first-order Laplacian matrices of the spatial and spectral graphs are defined as
@@ III-B Multi-Order Graph
Although constructing first-order nearest neighbor graphs can achieve a low-dimensional approximation of adjacent pixel representations [63], such local connectivity constraints often fail to characterize the global information inherent in HSIs, where the similarity between many pixels may extend beyond direct neighbors. In this regard, high-order graphs can be considered to improve the accuracy of HU [64]. Let $\mathbf{W}\in\mathbb{R}^{N\times N}$ denote the first-order graph with $N$ nodes, and the $K$-order graph $\mathbf{W}_{K}$ is defined as
@@
$\mathbf{W}_{m}$ $=\varrho_{1}\mathbf{W}_{1}+\varrho_{2}\mathbf{W}_{2}+\dots+\varrho_{K}\mathbf{W}_{K},$ (12)
$\mathbf{D}_{m}$ $=\varrho_{1}\mathbf{D}_{1}+\varrho_{2}\mathbf{D}_{2}+\dots+\varrho_{K}\mathbf{D}_{K},$
$\mathbf{L}_{m}$ $=\varrho_{1}\mathbf{L}_{1}+\varrho_{2}\mathbf{L}_{2}+\dots+\varrho_{K}\mathbf{L}_{K},$
where $\varrho_{1},\dots,\varrho_{k}$ are the weight parameters. From both spatial and spectral perspectives, we construct the following spatial and spectral multi-order graph
$\mathbf{W}_{m}$ $=\varrho_{1}\mathbf{W}_{1}^{\text{spa}}+\varrho_{2}\mathbf{W}_{2}^{\text{spa}}+\dots+\varrho_{K}\mathbf{W}_{K}^{\text{spa}}$ (13)
@@
$\mathbf{D}_{m}$ $=\varrho_{1}\mathbf{D}_{1}^{\text{spa}}+\varrho_{2}\mathbf{D}_{2}^{\text{spa}}+\dots+\varrho_{K}\mathbf{D}_{K}^{\text{spa}}$
@@
$\mathbf{L}_{m}$ $=\varrho_{1}\mathbf{L}_{1}^{\text{spa}}+\varrho_{2}\mathbf{L}_{2}^{\text{spa}}+\dots+\varrho_{K}\mathbf{L}_{K}^{\text{spa}}$
@@
$=\text{Tr}(\mathbf{S}\mathbf{L}_{m}\mathbf{S}^{\top}),$
where $\mathbf{L}_{m}$ is the adaptive Laplacian matrix. Compared to traditional $k$-NN or heat kernel-based graph constructions, this adaptive Laplacian matrix can be dynamically adjusted according to different hyperspectral data, thereby capturing local and global information more comprehensively. In addition, it can learn different order graphs in the spectral and spatial domains, which helps to filter out irrelevant connections caused by noise or sensor defects, making the graph structure more stable.
@@ III-C Our Model
s.t. $\mathbf{A}\geq 0,~\mathbf{S}\geq 0,~\mathbf{1}_{M}^{\top}\mathbf{S}=\mathbf{1}_{N}^{\top},$
@@ IV-A3 Update $\mathbf{E}$
Finally, Algorithm 1 summarizes the overall framework for solving (15), where the parameters of multi-order graphs in updating $\mathbf{S}$ can be adaptively learned through Algorithm 2. Here, $\mathcal{L}_{1}^{(i)}$ denotes the objective value after the $i$-th update.
@@
Initialize: $(\mathbf{A}^{0},\mathbf{S}^{0},\mathbf{E}^{0}),\;\epsilon_{1}=10^{-4},\;T_{1}=3000$
Repeat
@@
4: Compute objective value
$\mathcal{L}_{1}^{(i)}=\|\mathbf{X}-\mathbf{A}^{(i)}\mathbf{S}^{(i)}\|_{F}^{2}$
5: Set $i\leftarrow i+1$
Until $|\mathcal{L}_{1}^{(i)}-\mathcal{L}_{1}^{(i-1)}|<\epsilon_{1}$ or $i\geq T_{1}$
@@ IV-B1 Update $\mathbf{W}_{m}$
$\mathbf{W}_{m}\leftarrow\max\left(0,\sum_{v=1}^{V}\sum_{k=1}^{K}H_{vk}\mathbf{W}_{k}^{v}/\left(1+\mu\right)\right).$ (30)
@@ IV-B2 Update $\mathbf{H}$
Initialize: $H_{vk}=\frac{1}{2K},\epsilon_{2}=10^{-6},\;T_{2}=50$, multi-order graphs via (10)
Repeat
@@
3: Compute objective value:
$\mathcal{L}_{2}^{(j)}=\sum_{v=1}^{V}\sum_{k=1}^{K}H_{vk}^{(j)}\|\mathbf{W}_{m}^{(j)}-\mathbf{W}_{k}^{v}\|_{F}^{2}+\mu\|\mathbf{W}_{m}^{(j)}\|_{F}^{2}+\alpha\|\mathbf{H}^{(j)}\|_{F}^{2}$
4: Set $j\leftarrow j+1$
Until $|\mathcal{L}_{2}^{(j)}-\mathcal{L}_{2}^{(j-1)}|<\epsilon_{2}$ or $j\geq T_{2}$
@@
Remark IV.1
We would like to point out that: (1) For the $\mathbf{S}$-subproblem, $\ell_{1/2}$-norm is introduced to promote sparsity in the abundance maps, but it is nonsmooth and nonconvex. (2) For the $\mathbf{E}$-subproblem, $\ell_{2,1}$-norm is incorporated to enhance robustness against noise and outliers, but it is nonsmooth yet convex. (3) In the experimental part, the proposed algorithm is validated to exhibit a monotonically decreasing objective value during the optimization process.
IV-C Computational Complexity Analysis
IV-C1 Time Complexity
In Algorithm 1, the computational complexity of standard NMF is $\mathcal{O}(LNM)$. The update of $\mathbf{E}$ involves applying soft-thresholding to each row, with a complexity of $\mathcal{O}(LN)$. In addition, updating $\mathbf{S}$ requires computing $\mathbf{S}^{-1/2}$, which incurs a computational cost of $\mathcal{O}(MN)$. The computation of $\mathbf{S}\mathbf{W}_{m}$ and $\mathbf{S}\mathbf{D}_{m}$ both incur a complexity of $\mathcal{O}(MN^{2})$. Therefore, the overall computational complexity of Algorithm 1 is given by
$\mathcal{O}($ $LNM+LN+MN+MN^{2}).$ (33)
In Algorithm 2, the computational complexity of updating $\mathbf{W}_{m}$ is $\mathcal{O}(2KN^{2})$. Updating $\mathbf{H}$ involves solving a QP problem, whose complexity is $\mathcal{O}((VK)^{3})$. Since $V$ and $K$, representing the number of views and the graph order respectively, are relatively small, the cost of updating $\mathbf{H}$ can be considered negligible. It is worth noting that the algorithm requires constructing first-order graphs from both spectral and spatial views. We use nearest neighbour $C$, and the corresponding C-nearest neighbor graphs have a computational cost of $\mathcal{O}(N^{2}L)$ and $\mathcal{O}(N^{2})$, respectively. For each higher-order graph, the second-order weight matrix incurs a cost of $\mathcal{O}(C^{2}N)$. Therefore, by recursive reasoning, the total complexity for constructing $K$-order graphs can be derived as
$\sum_{i=2}^{K}\mathcal{O}(C^{2}N)=(K-1)\cdot\mathcal{O}(C^{2}N)=\mathcal{O}(KC^{2}N).$ (34)
Therefore, the overall computational complexity of Algorithm 2 is given by
$\mathcal{O}(2KN^{2}+N^{2}L+N^{2}+KC^{2}N).$ (35)
If Algorithm 1 terminates after $\psi$ iterations, and Algorithm 2 terminates after $\upsilon$ iterations, then the total computational cost can be expressed as
$\mathcal{O}($ $\psi(LNM+LN+MN)+\upsilon(2KN^{2})$ (36)
$+N^{2}L+N^{2}+KC^{2}N+MN^{2}).$
IV-C2 Space Complexity
The space complexity of MOGNMF is determined by the memory usage of the data matrix $\mathbf{X}\in\mathbb{R}^{L\times N}$, the endmember matrix $\mathbf{A}\in\mathbb{R}^{L\times M}$, the abundance matrix $\mathbf{S}\in\mathbb{R}^{M\times N}$, the noise matrix $\mathbf{E}\in\mathbb{R}^{L\times N}$, the fused graph matrix $\mathbf{W}_{m}\in\mathbb{R}^{N\times N}$, and the multi-order weight matrix $\mathbf{H}\in\mathbb{R}^{2K\times 2K}$. Therefore, the overall space complexity is given by $\mathcal{O}(LN+LM+MN+N^{2}+4K^{2}).$ Since $K\ll N$, the term $4K^{2}$ can be considered negligible. Thus, the overall space complexity simplifies to
$\mathcal{O}(LN+LM+MN+N^{2}).$ (37)
@@ V Numerical Experiments
This section presents numerical experiments between our method and some state-of-the-art HU methods, including VCA-FCLS [^1] https://github.com/isaacgerg/matlabHyperspectralToolbox [30] [31], MVC-NMF [^2] https://github.com/aicip/MVCNMF [35], SNMF [^3] https://github.com/yuehniu/Remote.Sensing [36] , RNMF [^4] https://www.irit.fr/~Cedric.Fevotte/extras/tip2015/code.zip[47] , SGSNMF [^5] https://github.com/YW81/TGRS17-SGSNMF [40], IRNMF [48], HGNMF-FS [55], MV-DUAL [^6] https://github.com/mabdolali/MaxVol_Dual/ [49], and A2SAN [^7] https://github.com/xuanwentao/A2SN-and-A2SAN [58].
@@ V-A1 Parameter Settings
For our proposed MOGNMF, $\gamma$, $\beta$, and $\lambda$ should be taken into account. According to [36], $\gamma$ is estimated by
@@
where $L$ represents the number of spectral bands, $N$ represents the total number of pixels, and $\mathbf{x}^{l}$ represents the vector corresponding to the $l$-th band in the hyperspectral data $\mathbf{X}\in\mathbb{R}^{L\times N}$. The regularization parameters are selected from the candidate set $\{10^{-3},10^{-2},10^{-1},10^{0},10,10^{2},10^{3}\}$. For other compared methods, either the default values or the best parameters recommended by the original authors are adopted.